OpenStudy (anonymous):

How do you solve 0=sin^2 x

OpenStudy (anonymous):

@SolomonZelman can you help me? :)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 0=\sin^2 x }\) \(\normalsize\color{blue}{ 0=1-\cos^2 x }\) \(\normalsize\color{blue}{ \cos^2 x=1 }\) \(\normalsize\color{blue}{ \cos x= ± 1 }\) but `│ cos θ│ < 1` so no solution

OpenStudy (solomonzelman):

`│ cos θ│ < 1` because `cos=Adjacent/Hypotenuse` and `Hyp>Adj`

OpenStudy (anonymous):

i forgot to mention this but the solutions of the equation have to be in the interval (0, 2pi). would there be a solution now?

OpenStudy (solomonzelman):

You can also know that there is no solution, becuase \(\normalsize\color{blue}{ \sin^2x=0 }\) gives, \(\normalsize\color{blue}{ \sin x=0 }\) and that gives, \(\normalsize\color{blue}{ 0= \rm { adjacent ~/~ hypotenuse} }\) but a side can not be equal to zero.

OpenStudy (solomonzelman):

I am just thinking abstractly.

OpenStudy (anonymous):

ok thanks!